# Inflated hot Jupiters: inferring average atmospheric velocity via Ohmic models coupled with internal dynamo evolution

### v1

@@ 6 Discussion and conclusions
In this work we have investigated the evolution of Jupiter-like planets, subject to both irradiation and internal Ohmic heating, via a suite of numerical simulations with the code MESA. We adopt time-dependent radial profiles of the electrical conductivity $\sigma(p,T)$, obtained by interpolations of the latest calculations for thermal alkali ionization (Kumar et al., 2021) and the degenerate metallic hydrogen contribution evaluated via DFT-MD calculations (Bonitz et al., 2024). This refines further the profiles of conductivity presented e.g. by Huang & Cumming (2012); Wu & Lithwick (2013), who accounted for pressure ionization but neglected strong coupling effects. Compared to those studies, our model yields conductivities in the metallic hydrogen region that are higher by approximately one order of magnitude, and in line with French et al. (2012). Although the precise value of $\sigma$ in the deep interior is largely irrelevant, since most of the deposited heat is deposited in much more shallower regions, we have included these values for completeness, and to enable a more accurate assessment of the intensity of the induced current intensity.
@@
Given that in the wind region the problem is intrinsically 3D, in our 1D MESA-adapted code we parametrize the atmospheric currents as proportional to the time-varying background magnetic field, estimated through usual dynamo scaling laws, the average conductivity in the uppermost layers, and a free parameter, $v_{\rm avg}$, which represent the average of the thermal winds in the outermost layers (eq. 10). In order to prescribe the profile of the current density at larger depths, we follow the classical approach by Batygin & Stevenson (2010) and Batygin et al. (2011) to find induction equation solutions, and we propose new semi-analytical solutions for the internally induced currents, based on piece-wise power-law fits of any conductivity profiles, and on any multipolar combination of the atmospheric wind flows and magnetic field geometry. The radial profiles of induced currents depend mostly on the radial dependence of the conductivity, which is steep except in the innermost region. The wind and magnetic geometries play a much lesser role, as also mentioned for a few specific simple cases in Batygin et al. (2011).
@@
With our model, we explore the combined dependence of the planetary radius evolution on the equilibrium temperature, the planetary mass, and the average velocity $v_{\rm avg}$ (for a given core mass and composition). As the planet cools down, both the dynamo magnetic field, $B_{\rm dyn}(t)$, and the temperature-controlled atmospheric conductivity decrease, with the net effect of a relevant decrease of the Ohmic efficiency in the long-term. Therefore, also the planetary radii decrease in time, and it is hard to match large values of $R\sim 1.5\leavevmode\nobreak\ R_{j}$ for moderate irradiation $T_{\rm eq}\sim 1500$ K.
This contrasts with previous studies that adopt the Ohmic heating scenario (e.g., Batygin & Stevenson 2010; Batygin et al. 2011; Knierim et al. 2022) or those that leave the underlying mechanism unspecified (e.g., Komacek & Youdin 2017; Thorngren & Fortney 2018), in which the heating efficiency is assumed to remain constant over time, and an equilibrium between the cooling and heating rates is reached, bringing the internal structure to be basically unchanged for many Gyr.
@@
Secondly, the amount of radius inflation is strongly dependent on the parameter $v_{\rm avg}$, and we have evaluated the range of values needed to explain the bulk of measured radii for HJs with a given $T_{\rm eq}$. The inferred decrease of $v_{\rm avg}$ with $T_{\rm eq}$ is much steeper than the one with $M$. Typically, we need $v_{\rm avg}$ of km/s in the lightest, moderately irradiated planets, and one or two orders of magnitude less when we move up to the highest values of $M$ and $T_{\rm eq}$ here considered. Keeping in mind that $v_{\rm avg}$ is a proxy to the wind velocity averaged over the uppermost region, $p\lesssim 10$ bar, these average values are compatible with GCMs, which find maximum velocities of a few km/s. It is also compatible with the study by Koll & Komacek (2018).
@@
The decrease of $v_{\rm avg}$ with $M$ is a consequence of the dependence on mass of the estimated internal field $B_{\rm dyn}$, which sets the atmospheric background field $B_{\rm atm}$, to which the induced currents are proportional, $J_{atm}\propto\sigma_{\rm atm}B_{\rm dyn}v_{\rm avg}$. Since there is a lack of highly inflated planets for high masses (Sestovic et al., 2018; Thorngren, 2024), this points to low average velocities for very massive planets. Both this fact and the decrease of $v_{\rm avg}$ with $T_{\rm eq}$ are consistent with the expected effects of the magnetic drag in slowing down the winds in the high temperature regime. Past works (Menou, 2012a; Knierim et al., 2022) also pointed to a steep decrease of wind velocities with $T_{\rm eq}$. Although 1D evolutionary models simulate the bulk structure of the planet and therefore cannot directly predict wind velocities at pressures $\lesssim$ mbar as probed by high-resolution spectroscopy, the trends of the average wind velocities with planetary mass and $T_{\rm eq}$ can still be tested with a sufficiently large sample.
@@
Another interesting point is that the typical induced current densities that penetrate the dynamo region are of the same order as the ones sustaining the deep-seated field, which we have estimated quantitatively by a combination of dynamo simulations and the typical magnetospheric field line curvature extracted from Juno data. This casts the question on how the atmospheric induced currents are going to couple with the ones sustaining the dynamo in the deep regions, and calls for dynamo simulations with different boundary conditions (usually taken as current-free and designed for cold giants, e.g. Duarte et al. 2013, 2018; Wicht et al. 2019b, a; Elias-López et al. 2025b).
@@
In our models there is a strong interplay between the value of the background field, generated via dynamo in the metallic hydrogen region, and the Ohmic efficiency, due to the proportionality between the induced and background fields. An important effect is that, for strongly heated models, the convection region can be strongly suppressed, as noted also by Elias-López et al. (2025a). The drop in the Ohmic efficiency allows the convection region to be recovered, giving cycles of intermittent dynamo and induced field. This feedback effect has been neglected in previous studies, which assume constant values for both the Ohmic efficiency and the background field, with no feedback effects or use of convective heat scaling laws.
@@
The specifics of this effect, and an in-depth understanding of the characteristic timescales over which these cycles may operate, cannot be properly addressed by this study. In fact, the frequency and timing of these cycles are rather sensitive to both the physical and numerical choices made here, such as the evolution time step, the grid resolution, the form of the scaling laws employed (see App. A), and the temporal average applied to the heat source term for numerical reasons (Sect. 4.4). However, these numerical factors alone are not the ones responsible for the occurrence of such behaviour, which arises regardless of these details (albeit with variation in form), and has a solid physical explanation: the circular relation between induced Ohmic dissipation, convective heat flux and dynamo magnetic field. A more quantitative study of this behaviour would require to address the problem considering significantly smaller timescales than the ones used here. Although MESA automatically adapts them close to the cycles, they can become so short to potentially compete with other more dynamical timescales, neglected by the long-term cooling code, and related for instance to convection or dynamo magnetic cycles. In other words, variations over these timescales would require to consider effects which are typically neglected in a long-term evolutionary model, which implicitly averages out oscillatory effects acting on short timescales. Moreover, this effect relies on the assumption that the average velocity is constant and is not affected by a change in the magnetic field, and on magnetic scaling laws which also implicitly provide a simple estimate of the average dynamo field, not considering short-term fluctuations.
@@
Related to the last point, this feedback effect would be probably mitigated by the magnetic drag, for which the wind velocity will also quickly respond to compensate any important decrease or increase of the background temperature, something we do not consider (we assume the average velocity to remain constant over time). Another possible caveat that could have enhanced the effects of the coupling in this study is the linear dependence of the induced currents with the deep-seated field: when the induced magnetic field becomes comparable with the latter (magnetic Reynolds number $Rm\gg 1$), which is expected for a large fraction of HJs (Batygin et al., 2013; Dietrich et al., 2022; Soriano-Guerrero et al., 2025), the induction mechanism is not linear, and the assumption of linear proportionality $J\propto B_{\rm dyn}$ is likely simplistic.
@@
We have also briefly explored the effects of an evolving stellar luminosity, rather than assuming it constant. While the main effects discussed here (coupling between dynamo-generated field and induced currents, overall trends, change in Ohmic efficiency) remain largely valid, our results show re-inflation at late stages (several Gyr), in agreement with previous studies (Lopez & Fortney, 2016; Komacek et al., 2020), and with observational hints (Hartman et al., 2016; Thorngren et al., 2021). We have restricted our analysis to the main-sequence stage of a Sun-like star, leaving a more extensive parameter exploration to future studies. The evolving stellar irradiation further underscores the need to move beyond the simplifying assumption of a constant $v_{\rm avg}$.
@@
In this sense, future works need to consider with further detail the interplay between the deep-seated dynamo and the atmospheric induction with more detail. Another ingredient to be potentially taken into account in the future is the double-diffusive convection, which might become relevant when the convection is suppressed in highly heated HJs. In any case, our current findings have potential consequences on the elusive radio detectability of HJs (Stevens, 2005; Grießmeier et al., 2007; Zarka, 2007), which, so far, has only given, in the best case, unconfirmed claims (see e.g. Turner et al. 2021, 2024), despite multiple observational campaigns and promising theoretical predictions. Finally, the inferred values of the average velocity as a function of $M$ and $T_{\rm eq}$ can be put in the context of population synthesis (Thorngren & Fortney, 2018), considering also other variables (core size and composition, atmospheric modelling), here taken as constant, which might introduce dispersion in the observed radii.
@@ Appendix A On the different scaling laws
Some words of caution must be spent about the applicability of the dynamo field scaling laws, used here and in several works. The Christensen et al. (2009) scaling law employed in this study, eq. (18), has seen to hold in a broad range of masses, from Earth and Jupiter to main-sequence stars, with a $\sim 60\%$ scattering in the proportionality pre-coefficient of the scaling laws (Reiners et al. 2009). However, notice that several brown dwarfs are outliers in both directions: surprisingly low magnetic field for infant, accreting ones (Reiners et al. 2009), and high (kG) values for rapidly rotating ones, as inferred from coherent, polarized radio emission at GHz (Kao et al. 2016, 2018). Moreover, Christensen et al. (2009) explicitly discarded Mercury, Saturn, Uranus and Neptune from the scaling laws, substantially due to the likely absence of a thick convective region. Therefore, despite the remarkable range of heat flux and magnetic field range covered by the correlation, the planetary application is inferred to be non-trivial in the original work, and needs further considerations.
@@
Figure 10: Magnetic field estimates, for a $M=1\leavevmode\nobreak\ M_{j}$ planet, with $T_{\rm eq}=300$ K (a cold Jupiter, cyan), 1500 K (black), and 2000 K (grey), and for $T_{\rm eq}=1500$ K plus Ohmic heating with $v_{\rm avg}=3$ km/s, but fixing $B_{\rm atm}=10$ G (red, see text for the motivation). The different line styles correspond to the estimates by Christensen et al. (2009) (thick lines), eq. (18) with $f_{\rm ohm}=0.5$, using the integral in the nominal dynamo region only ($p>p_{\rm dyn}=10^{6}$ bar, solid lines) or in the entire convective region (dashes), and by Reiners et al. (2009), eq. (22) (thin lines), using the internal luminosity at $p=10^{6}$ bar (solid) or at $p=10^{3}$ bar (dashed).
@@
In Fig. 10 we show the different magnetic field estimates $B_{\rm dyn}$, for a $M=1\leavevmode\nobreak\ M_{j}$ planet. We consider four cases: $T_{\rm eq}=300,1500,2000$ K without additional heating ($\epsilon_{\rm j}=0$), and for $T_{\rm eq}=1500$ K plus Ohmic heating. The latter is run with $v_{\rm avg}=3$ km/s, but fixing $B_{\rm atm}=10$ G in the $J$ normalization, eq. (10), instead of evolving it with the same estimate of $B_{\rm dyn}$ via eq. (19), in order to decouple the differences in the estimates from the feedback of $B_{\rm dyn}$ on the induced $J$ and on the cooling, discussed in the main text.
@@
Several things must be noticed. First of all, large differences are seen at very early ages, which is related to the discussion in Sect. 4.4. More importantly, at times relevant for this study, $t\gtrsim 10^{8}$ yr, the plot shows that the formula by Christensen et al. (2009) is always pretty insensitive to the minimum pressure from which the integral in eq. (18) is evaluated. This is due to the convective heat flux $q_{c}(r)$, which is zero in the radiative region, and, together with the density, gives the bulk of the contribution to the integral from the dynamo region. Secondly, the simplified formula (22) evaluated at $p_{\rm dyn}$ is seen to systematically provide a factor $\sim 3$ less than the original one, in a broad range of planetary models (including different planetary masses, $T_{\rm eq}$ and $v_{\rm avg}$, here not shown). The two curves would be compatible between each other for a smaller choice of $f_{\rm dip}\sim 0.06$ (but notice that in the original work, Christensen et al. (2009) employed $f_{\rm ohm}=1$, and numerical simulations typically provide $f_{\rm ohm}\sim 0.3-0.7$, Christensen & Aubert 2006).
@@
Besides this, the most important result is that, in case of a heated HJ (red curves), the choice of eq. (22) evaluated at $p=10^{3}$ bar, provides a radically different behaviour, being almost constant in time after the Ohmic term is turned on. This is due to the fact that the additionally heat in the outer convective region, with time, becomes the most relevant contribution and provides a constant-in-time internal luminosity to balance the radiated energy. Note that Yadav & Thorngren (2017) use the deposited heat as an input value for the scaling law, and they don’t specify neither the mechanism nor its specific heating profile. If their approach is applied to the Ohmic dissipation, it presents conceptual issues: (i) the extra luminosity originates from the dissipation of the atmospherically-induced field in the shallow layers, while the heat flux in the Christensen et al. (2009) scaling law has to be evaluated in the deep-seated, highly conductive dynamo region alone; (ii) having Ohmic luminosity both as a dissipation of magnetic field and a source of it via convective flux (even if in different regions) is unclear in terms of energy conservation. In our models, which don’t fix $B_{\rm atm}$, the use of eq. (22) evaluated at $p=10^{3}$ bar would provide a more complicated curve, due to the coupling, eqs. (10) and (19), but the arguments above still hold.
@@
Finally, keep in mind an additional caveat which applies to any version of the scaling law. As discussed in the main text and shown by previous works (Komacek & Youdin 2017; Komacek et al. 2020; Elias-López et al. 2025a) a strong deposition of heat in Ohmic models (with the heat concentrated in the outer region) can lead to a fragmentation of the large-scale convection into thinner layer(s). This kind of dynamo deviates from the almost fully convective objects against which the scaling laws were calibrated (Christensen et al. 2009), in a similar way to the cases of the aforementioned majority of Solar planets with thin dynamos, excluded a-priori from the original correlation.

### v2

@@ 6 Discussion and conclusions
In this work we have investigated the evolution of Jupiter-like planets subject to both irradiation and internal Ohmic heating via a suite of numerical simulations with the code MESA. We adopted time-dependent radial profiles of the electrical conductivity $\sigma(p,T)$ obtained by interpolations of the latest calculations for thermal alkali ionisation (Kumar et al., 2021) and the degenerate metallic hydrogen contribution evaluated via DFT-MD calculations (Bonitz et al., 2024). Employing these values further refines the profiles of conductivity presented, for example, by Huang & Cumming (2012); Wu & Lithwick (2013), who accounted for pressure ionisation but neglected strong coupling effects. Compared to those studies, our model yields conductivities in the metallic hydrogen region that are higher by approximately one order of magnitude and in line with French et al. (2012). Although the precise value of $\sigma$ in the deep interior is largely irrelevant, since most of the deposited heat is deposited in much more shallower regions, we have included these values for completeness and to enable a more accurate assessment of the intensity of the induced current intensity.
@@
Given that the problem in the wind region is intrinsically 3D, in our 1D MESA-adapted code we parametrised the atmospheric currents as proportional to the time-varying background magnetic field, estimated through usual dynamo scaling laws, the average conductivity in the uppermost layers, and a free parameter, $v_{\rm avg}$, which represents the average of the thermal winds in the outermost layers (eq. 10). To prescribe the profile of the current density at larger depths, we followed the classical approach by Batygin & Stevenson (2010) and Batygin et al. (2011) to find induction equation solutions, and we have proposed new semi-analytical solutions for the internally induced currents based on piece-wise power-law fits of any conductivity profile and on any multipolar combination of the atmospheric wind flows and magnetic field geometry. The radial profiles of induced currents depend mostly on the radial dependence of the conductivity, which is steep except in the innermost region. The wind and magnetic geometries play a much less significant role, as also mentioned for a few specific simple cases in Batygin et al. (2011).
@@
With our model, we explored the combined dependence of the planetary radius evolution on the equilibrium temperature, the planetary mass, and the average velocity, $v_{\rm avg}$ (for a given core mass and composition). As the planet cools down, both the dynamo magnetic field, $B_{\rm dyn}(t)$, and the temperature-controlled atmospheric conductivity decrease, with the net effect of a relevant decrease of the Ohmic efficiency in the long term. Therefore, also the planetary radii decrease in time, and it is hard to match large values of $R\sim 1.5\leavevmode\nobreak\ R_{\rm j}$ for moderate irradiation $T_{\rm eq}\sim 1500$ K. This contrasts with previous studies that adopted the Ohmic heating scenario (e.g. Batygin & Stevenson 2010; Batygin et al. 2011; Knierim et al. 2022) or those that left the underlying mechanism unspecified (e.g. Komacek & Youdin 2017; Thorngren & Fortney 2018), in which case the heating efficiency was assumed to remain constant over time. In those cases, an equilibrium between the cooling and heating rates is reached, bringing the internal structure to be basically unchanged for many gigayears.
@@
Secondly, the amount of radius inflation is strongly dependent on the parameter $v_{\rm avg}$, and we evaluated the range of values needed to explain the bulk of measured radii for HJs with a given $T_{\rm eq}$. The inferred decrease of $v_{\rm avg}$ with $T_{\rm eq}$ is much steeper than the one with $M$. Typically, we need $v_{\rm avg}\sim$ km/s in the lightest, moderately irradiated planets, and one or two orders of magnitude less when we move up to the highest values of $M$ and $T_{\rm eq}$ considered here. Keeping in mind that $v_{\rm avg}$ is a proxy of the wind velocity averaged over the uppermost region, $p\lesssim 10$ bar, these average values are compatible with GCMs, which find maximum velocities of a few kilometers per second. It is also compatible with the study by Koll & Komacek (2018).
@@
The decrease of $v_{\rm avg}$ with $M$ is a consequence of the dependence on the mass of the estimated internal field, $B_{\rm dyn}$, which sets the atmospheric background field, $B_{\rm atm}$, which in turn enters in the normalization of the induced currents, $J_{\rm atm}\propto\sigma_{\rm atm}B_{\rm dyn}v_{\rm avg}$. Since there is a lack of highly inflated planets for high masses (Sestovic et al., 2018; Thorngren, 2024), this points to low average velocities for very massive planets. Both this fact and the decrease of $v_{\rm avg}$ with $T_{\rm eq}$ are consistent with the expected effects of the magnetic drag in slowing down the winds in the high temperature regime. Past works (Menou, 2012a; Knierim et al., 2022) have also pointed to a steep decrease of wind velocities with $T_{\rm eq}$. Although 1D evolutionary models simulate the bulk structure of the planet and therefore cannot directly predict wind velocities at pressures $\lesssim$ mbar as probed by high-resolution spectroscopy, the trends of the average wind velocities with planetary mass and $T_{\rm eq}$ can still be tested with a sufficiently large sample.
@@
Another interesting point is that the typical induced current densities that penetrate the dynamo region are of the same order as the ones sustaining the deep-seated field, which we have estimated quantitatively by a combination of dynamo simulations and the typical magnetospheric field line curvature extracted from Juno data. This raises the question of how the atmospheric induced currents are going to couple with the ones sustaining the dynamo in the deep regions and calls for dynamo simulations with different boundary conditions (which are usually taken as current-free and designed for cold giants, e.g. Duarte et al. 2013, 2018; Wicht et al. 2019b, a; Elias-López et al. 2025b).
@@
In our models there is a strong interplay between the value of the background field, generated via a dynamo in the metallic hydrogen region, and the Ohmic efficiency, due to the proportionality between the induced and background fields. An important effect is that for strongly heated models, the convection region can be strongly suppressed, as noted also by Elias-López et al. (2025a). The drop in the Ohmic efficiency allows the convection region to be recovered, giving cycles of intermittent dynamo and induced field. This feedback effect has been neglected in previous studies, which assume constant values for both the Ohmic efficiency and the background field, with no feedback effects or use of convective heat scaling laws.
@@
The specifics of this effect and an in-depth understanding of the characteristic timescales over which these cycles may operate cannot be properly addressed by this study. In fact, the frequency and timing of these cycles are rather sensitive to both the physical and numerical choices made here, such as the evolution time step, the grid resolution, the form of the scaling laws employed (see App. A), and the temporal average applied to the heat source term for numerical reasons (Sect. 4.4). However, these numerical factors alone are not the ones responsible for the occurrence of such behaviour, which arises regardless of these details (albeit with a variation in form), and has a solid physical explanation: the circular relation between induced Ohmic dissipation, convective heat flux, and dynamo magnetic field. A more quantitative study of this behaviour would require addressing the problem of considering significantly smaller timescales than the ones used here. Although MESA automatically adapts them close to the cycles, they can become so short as to potentially compete with other more dynamical timescales neglected by the long-term cooling code and related, for instance, to convection or dynamo magnetic cycles. In other words, variations over these timescales would require consideration of the effects that are typically neglected in a long-term evolutionary model, which implicitly averages out oscillatory effects acting on short timescales. Moreover, this effect relies on the assumption that the average velocity is constant and is not affected by a change in the magnetic field and on magnetic scaling laws, which also implicitly provide a simple estimate of the average dynamo field, not considering short-term fluctuations.
@@
Related to the last point, this feedback effect would probably be mitigated by the magnetic drag, for which the wind velocity will also quickly respond to compensate for any important decrease or increase of the background temperature, something we do not consider (we assumed the average velocity remains constant over time). Another possible caveat that could have enhanced the effects of the coupling in this study is the linear dependence of the induced currents with the deep-seated field: When the induced magnetic field becomes comparable with the latter (magnetic Reynolds number $Rm\gg 1$), which is expected for a large fraction of HJs (Batygin et al., 2013; Dietrich et al., 2022; Soriano-Guerrero et al., 2025), the induction mechanism is not linear, and the assumption of linear proportionality, $J\propto B_{\rm dyn}$, is likely simplistic.
@@
We also briefly explored the effects of an evolving stellar luminosity rather than assuming it to be constant. While the main effects discussed here (coupling between dynamo-generated field and induced currents, overall trends, change in Ohmic efficiency) remain largely valid, our results show re-inflation at late stages (several gigayears), in agreement with previous studies (Lopez & Fortney, 2016; Komacek et al., 2020), and with observational hints (Hartman et al., 2016; Thorngren et al., 2021). We restricted our analysis to the main-sequence stage of a Sun-like star, leaving a more extensive parameter exploration to future studies. The evolving stellar irradiation further underscores the need to move beyond the simplifying assumption of a constant $v_{\rm avg}$.
@@
In this sense, future works need to consider with further detail the interplay between the deep-seated dynamo and the atmospheric induction with more detail. Another ingredient to potentially be taken into account in the future is the double-diffusive convection, which might become relevant when the convection is suppressed in highly heated HJs. In any case, our current findings have potential consequences on the elusive radio detectability of HJs (Stevens, 2005; Grießmeier et al., 2007; Zarka, 2007), which so far has only given, in the best case, unconfirmed claims (see e.g. Turner et al. 2021, 2024) despite multiple observational campaigns and promising theoretical predictions. Finally, the inferred values of the average velocity as a function of $M$ and $T_{\rm eq}$ can be put in the context of population synthesis (Thorngren & Fortney, 2018), considering also other variables (core size and composition, atmospheric modelling), which we took to be constant in this work and might introduce dispersion in the observed radii.
@@ Appendix A On the different scaling laws
Some words of caution must be spent about the applicability of the dynamo field scaling laws, used here and in several works. The Christensen et al. (2009) scaling law employed in this study, eq. (18), has seen to hold in a broad range of masses, from Earth and Jupiter to main-sequence stars, with a $\sim 60\%$ scattering in the proportionality pre-coefficient of the scaling laws (Reiners et al. 2009). However, we noticed that several brown dwarfs are outliers in both directions: surprisingly low magnetic field for infant, accreting ones (Reiners et al. 2009), and high (kG) values for rapidly rotating ones, as inferred from coherent, polarised radio emission at GHz (Kao et al. 2016, 2018). Moreover, Christensen et al. (2009) explicitly discarded Mercury, Saturn, Uranus and Neptune from the scaling laws, substantially due to the likely absence of a thick convective region. Therefore, despite the remarkable range of heat flux and magnetic field range covered by the correlation, the planetary application is inferred to be non-trivial in the original work, and needs further considerations.
@@
Figure 10: Magnetic field estimates, for a $M=1\leavevmode\nobreak\ M_{\rm j}$ planet, with $T_{\rm eq}=300$ K (a cold Jupiter, cyan), 1500 K (black), and 2000 K (grey), and for $T_{\rm eq}=1500$ K plus Ohmic heating with $v_{\rm avg}=3$ km/s, but fixing $B_{\rm atm}=10$ G (red, see text for the motivation). The different line styles correspond to the estimates by Christensen et al. (2009) (thick lines), eq. (18) with $f_{\rm ohm}=0.5$, using the integral in the nominal dynamo region only ($p>p_{\rm dyn}=10^{6}$ bar, solid lines) or in the entire convective region (dashes), and by Reiners et al. (2009), eq. (22) (thin lines), using the internal luminosity at $p=10^{6}$ bar (solid) or at $p=10^{3}$ bar (dashed).
@@
In Fig. 10 we show the different magnetic field estimates $B_{\rm dyn}$, for a $M=1\leavevmode\nobreak\ M_{\rm j}$ planet. We consider four cases: $T_{\rm eq}=300,1500,2000$ K without additional heating ($\epsilon_{\rm j}=0$), and for $T_{\rm eq}=1500$ K plus Ohmic heating. The latter is run with $v_{\rm avg}=3$ km/s, but fixing $B_{\rm atm}=10$ G in the $J$ normalisation, eq. (10), instead of evolving it with the same estimate of $B_{\rm dyn}$ via eq. (19), in order to decouple the differences in the estimates from the feedback of $B_{\rm dyn}$ on the induced $J$ and on the cooling, discussed in the main text.
@@
Several things must be noticed. First of all, large differences are seen at very early ages, which is related to the discussion in Sect. 4.4. More importantly, at times relevant for this study, $t\gtrsim 10^{8}$ yr, the plot shows that the formula by Christensen et al. (2009) is always pretty insensitive to the minimum pressure from which the integral in eq. (18) is evaluated. This is due to the convective heat flux $q_{c}(r)$, which is zero in the radiative region, and, together with the density, gives the bulk of the contribution to the integral from the dynamo region. Secondly, the simplified formula (22) evaluated at $p_{\rm dyn}$ is seen to systematically provide a factor $\sim 3$ less than the original one, in a broad range of planetary models (including different planetary masses, $T_{\rm eq}$ and $v_{\rm avg}$, here not shown). The two curves would be compatible between each other for a smaller choice of $f_{\rm dip}\sim 0.06$ (but we noticed that in the original work, Christensen et al. (2009) employed $f_{\rm ohm}=1$, and numerical simulations typically provide $f_{\rm ohm}\sim 0.3-0.7$, Christensen & Aubert 2006).
@@
Besides this, the most important result is that, in case of a heated HJ (red curves), the choice of eq. (22) evaluated at $p=10^{3}$ bar, provides a radically different behaviour, being almost constant in time after the Ohmic term is turned on. This is due to the fact that the additionally heat in the outer convective region, with time, becomes the most relevant contribution and provides a constant-in-time internal luminosity to balance the radiated energy. We note that Yadav & Thorngren (2017) use the deposited heat as an input value for the scaling law, and they do not specify neither the mechanism nor its specific heating profile. If their approach is applied to the Ohmic dissipation, it presents conceptual issues: (i) the extra luminosity originates from the dissipation of the atmospherically induced field in the shallow layers, while the heat flux in the Christensen et al. (2009) scaling law has to be evaluated in the deep-seated, highly conductive dynamo region alone; (ii) having Ohmic luminosity both as a dissipation of magnetic field and a source of it via convective flux (even if in different regions) is unclear in terms of energy conservation. In our models, which do not fix $B_{\rm atm}$, the use of eq. (22) evaluated at $p=10^{3}$ bar would provide a more complicated curve, due to the coupling, eqs. (10) and (19), but the arguments above still hold.
@@
Finally, keep in mind an additional caveat which applies to any version of the scaling law. As discussed in the main text and shown by previous works (Komacek & Youdin 2017; Komacek et al. 2020; Elias-López et al. 2025a) a strong deposition of heat in Ohmic models (with the heat concentrated in the outer region) can lead to a fragmentation of the large-scale convection into thinner layer(s). This kind of dynamo deviates from the almost fully convective objects against which the scaling laws were calibrated (Christensen et al. 2009), in a similar way to the cases of the aforementioned majority of Solar planets with thin dynamos, excluded a priori from the original correlation.